Let {an}∞ n=1 be a convergent sequence with limit L such that B≤an forevery n≥N for some N∈ N and some B∈R. Prove that B≤L


1
Expert's answer
2021-02-11T11:40:10-0500

Let us prove by contraposition. Suppose that "B>L". Therefore, "B-L>0". Let "\epsilon =\frac{B-L}{2}". Since "\{a_n\}_{n=1}^{\infty}" be a convergent sequence with limit "L", there exist "M\in\mathbb N" such that "|a_n-L|<\epsilon" for every "n\ge M". Let "k\ge \max\{N, M\}". Then "-\epsilon< a_k-L<\epsilon". It follows that "a_k<\epsilon + L=\frac{B-L}{2}+L=\frac{B+L}{2}<\frac{B+B}{2}=B". Consequently, "a_k<B", and we have a contradiction with "a_k\ge B". Therefore, "B\le L."